### v1

# ISWC’24: International Semantic Web Conference, 2024, Baltimore, USA

Abstract
Hallucinations of large language models (LLMs) commonly occur in domain-specific downstream tasks, with no exception in ontology matching (OM). The prevalence of using LLMs for OM raises the need for benchmarks to better understand LLM hallucinations. The OAEI-LLM dataset is an extended version of the Ontology Alignment Evaluation Initiative (OAEI) datasets that evaluate LLM-specific hallucinations in OM tasks. We outline the methodology used in dataset construction and schema extension, and provide examples of potential use cases.
keywords:
[…]
1 Motivation
Large language models (LLMs) have shown incredible capabilities in natural language generation (NLG) and question answering (QA). In the Semantic Web community, LLMs have recently been used in ontology matching (OM). Although LLMs offer a strong background knowledge base for OM, they can sentiently generate a number of false mappings and miss some true mappings, resulting in lower precision or recall. This phenomenon has been observed in several recent papers in [1, 2, 3, 4, 5]. It is mainly caused by LLM hallucinations, where LLMs tend to generate synthesised answers when they do not have sufficient background knowledge or have biased domain knowledge.
For example, the conference track [6] is one of the Ontology Alignment Evaluation Initiative (OAEI) [7] campaigns that evaluates OM tasks in conference organisation. It includes several reference alignments between different ontologies related to research conference management. LLM hallucinations can cause an incorrect match of “http://cmt#Paper” and “http://confOf#Paper”. LLMs lack domain knowledge to infer that “paper” in conference submission (e.g. in the ConfOf ontology) may have a narrower intended meaning like “full-version paper” that is distinguished from other submission types such as “abstract” or “poster”. The correct match in this case is “http://cmt#PaperFullVersion” with “http://confOf#Paper”. LLMs may also experience hallucinations when they encounter a missing match of “http://cmt#Chairman” and “http://conference#Chair”. It is very difficult for LLMs to understand polysemous words, such as where the word “chair” refers to the person who chairs the conference, and not to the furniture used at the conference. These examples can lead to low-precision or low-recall issues: the former mismatches the entity “paper” with the entity “paper”, while the latter assumes that the entity “chairman” does not match the entity “chair”. It seems feasible to mitigate LLM hallucinations using fine-tuning, but there are few datasets that can be used to provide baseline references.
Established in 2004, OAEI has become the largest community of specified ontology matching challenges. Over the years, OAEI has developed several multi-ontology tracks across a wide range of topic domains. Almost all tracks use expert systems or machine learning-based systems to generate ground-truth reference mapping results. We observe that no tracks have covered the topic of LLM hallucinations in OM tasks. For this reason, we propose to extend the current OAEI datasets to construct a new dataset called OAEI-LLM, which serves to measure the degree of hallucination by LLMs in OM tasks. The new dataset compares the original human-labelled results with LLM-generated results, classifies different types of hallucinations made by different LLMs, and records the information with a new schema extension. We believe this dataset would benefit research in understanding LLM hallucinations in matching tasks and thereby research in improving LLM-driven OM.
2 Methodology
2.1 Dataset Construction
Figure 1 illustrates the procedure of constructing the dataset. The original OAEI datasets provide three files: the source ontology ($O_{s}$), the target ontology ($O_{t}$), and their OAEI reference ($R_{oaei}$). We implement an LLM-based OM system to generate the alignment file LLM-Alignment ($A_{llm}$). The system takes $O_{s}$ and $O_{t}$ as inputs and generates a set of predicted mappings. The matching assessment procedure detailed below will compare $R_{oaei}$ with $A_{llm}$ and find their differences.
Figure 1: Dataset Construction.
For each pair of mappings $(e1,e2)\in R_{oaei}$ and $(e1^{\prime},e2^{\prime})\in A_{llm}$, we consider that there are no LLM hallucinations if $(e1,e2)=(e1^{\prime},e2^{\prime})$. We only deal with LLM hallucinations that occur when $(e1,e2)\neq(e1^{\prime},e2^{\prime})$. These can be categorised into two different categories: (1) Missing: For a specific pair $(e1,e2)$, there does not exist any pair $(e1^{\prime},e2^{\prime})$ satisfies $(e1=e1^{\prime})\lor(e2=e2^{\prime})$. (2) Incorrect: For a specific pair $(e1,e2)$, there exists a pair $(e1^{\prime},e2^{\prime})$ satisfies $(e1=e1^{\prime}\land e2\neq e2^{\prime})\lor(e1\neq e1^{\prime}\land e2=e2^{\prime})$. For missing mappings, additional details on the mapping that occurs in $R_{oaei}$ or $A_{llm}$ are provided. For incorrect mappings, additional details on the misaligned entity and type of hallucinations are provided. The matching assessment determines the following types of hallucinations in incorrect mappings:
  • Align-up: LLMs map the entity to its intended superclass.
  • Align-down: LLMs map the entity into its intended subclass.
  • False-mapping: LLMs map the entity to an irrelevant entity.
  • Disputed-mapping: LLMs map the entity to a relevant entity, but it does not align with the OAEI reference. We call this mapping type “disputed” because we cannot guarantee the OAEI reference is always correct. LLMs may discover a more precise matching entity. OAEI-LLM also aims to use LLMs to help with evolving more precise reference alignments for OM tasks.
The matching assessment could be done via a semi-automatic process (e.g. human-in-the-loop evaluation) or a fully automated process (e.g. multiple LLM agent negotiation). The final results will be used to generate the OAEI-LLM dataset.
2.2 Schema Extension
The OAEI tracks use a general alignment format [8] derived from the Expressive and Declarative Ontology Alignment Language (EDOAL) [9]. A simple example to represent a valid match of “chairman” and “chair” is illustrated as follows: Each mapping is recorded in a tag <Cell>. The schema uses the tags <entity1> and <entity2> to record the pair of entities, the tag <measure> to present their similarity, and the tag <relation> to indicate their relations.
We extend the current mapping schema to record the new information related to LLM hallucinations. For each mapping, different types of hallucinations can occur in different LLMs. Each of them is recorded in the tag <hallucination>. In the extended version of the example (highlighted in blue), the first segment depicts the LLM hallucinations that occurred in Llama-3.1, which is a case of missing mapping on the LLM side. The second segment illustrates the LLM hallucinations that occurred in GPT-4o, indicating an incorrect mapping to the entity “conference chair”. The matching assessment classifies this mapping as “disputed mapping”.
<map>
<Cell>
<entity1 rdf:resource=’http://cmt#Chairman’/>
<entity2 rdf:resource=’http://conference#Chair’/>
<measure rdf:datatype=’xsd:float’>1.0</measure>
<relation>=</relation>
<hallucination>
<llm>llama-3.1</llm>
<category>missing</category>
<source>llm</source>
</hallucination>
<hallucination>
<llm>gpt-4o</llm>
<category>incorrect</category>
<entity rdf:resource=’http://conference#ConferenceChair’/>
<type>dispute-mapping</type>
</hallucination>
</Cell>
</map>
3 Potential Use Cases
3.1 Benchmarking LLMs for OM tasks
Different LLMs can achieve different performance on the same OM task. However, running an LLM-based matcher on a traditional OAEI dataset can only observe the differences in precision, recall, and F1 score. Using the extended version of OAEI-LLM, we can additionally quantify the LLM errors made by counting each of the different types of errors we have identified. Therefore, it becomes possible to understand the tendency of LLMs to generate incorrect answers or to provide relevant but not precise answers. This could also be a valuable supplement to enhance an LLM leaderboard for OM tasks. Indeed, this same additional information on matching failures could also be helpful if adopted more widely by non-LLM ontology matchers.
3.2 A dataset for fine-tuning LLMs used in OM tasks
Fine-tuning is a common approach used in LLMs to inject domain knowledge, but a significant amount of data typically needs to be provided. Unlike the traditional OAEI dataset, which only provides information on matched entities, OAEI-LLM provides extensive information identifying when and how LLM hallucinations occur, and even the mismatched entities caused by LLM hallucinations. This could be used to generate high-quality training data to fine-tune LLMs.
Ethical Considerations
According to the OAEI data policy (date accessed: 2024-06-30), “OAEI results and datasets, are publicly available, but subject to a use policy similar to the one defined by NIST for TREC. These rules apply to anyone using these data.” Please find more details from the official website: https://oaei.ontologymatching.org/doc/oaei-deontology.2.html
Acknowledgements.
The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.